OpenStudy (anonymous):

I know how to do this, but I'd like someone else to see if this answer is correct just for confirmation. the indefinite integral of (5x^5 + 1)^(2/3) * 125x^4 dx

OpenStudy (anonymous):

My answer was 12/5 (5x^5 + 1) ^ (5/3) + C

OpenStudy (michele_laino):

I got a different answer

OpenStudy (anonymous):

Okay let me walk through what I did

OpenStudy (michele_laino):

please try this substitution: \[5{x^5} + 1 = t\]

OpenStudy (anonymous):

I set u = to 5x^5 + 1 then related that to dx with du = 25x^4dx That left me with 4u^(2/3) dx I said the antiderivative of that was 12/5u^(5/3) <------ This may be the part I messed up on Then I filled everything back in for u and added C

OpenStudy (michele_laino):

no, it is 5*u^(2/3)

OpenStudy (anonymous):

Can you explain that? I'm awful at antiderivatives

OpenStudy (michele_laino):

since 125 x^4 dx= 5 * 25x^4 dx = 5 du

OpenStudy (michele_laino):

so your iontegrand function can be rewritten as ollows: 5 u^(2/3) du

OpenStudy (michele_laino):

and its antiderivative is: \[5\frac{{{u^{2/3 + 1}}}}{{\frac{2}{3} + 1}} + C\]

OpenStudy (anonymous):

Oh okay so my error was in saying 4 instead of 5?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Alright thank you so much. At least I'm getting the concept I think

OpenStudy (michele_laino):

thank you!